# Efficient Computation of Closed Substrings

### v1

@@ 4 Algorithms for Closed Substrings and MCSs
In this section, we introduce the primary data structure –the $\mathcal{MRC}$ array– which is used to compute all closed substrings and MCSs in $\mathcal{O}(n\log n)$ time.
Definition 1 ($\mathcal{MRC}$ Array)
@@
Given a string $w[1..n]$, the maximal right-closed array ($\mathcal{MRC}$) is an array of size $n$, where $\mathcal{MRC}[i]$, $1\leq i\leq n$, stores a list of ordered pairs $(\ell r_{j},\ell b_{j})$, where $j=K_{i}$ to $1$.
@@
Algorithm 1 computes the $\mathcal{MRC}$ array, by first identifying all the maximal right-closed substrings of length greater than one. It uses a stack to store AVL trees created during its execution. The algorithm scans the $\mathsf{LCP}$ array from left to right. Beginning with $\mathsf{LCP}[1]=0$, and for increasing $\mathsf{LCP}[i]$ values, the algorithm create a single node (root) AVL tree with $key=\mathsf{SA}[i]$ and $value=\mathsf{LCP}[i]$ and pushes it on the stack until either the $LCP$ value decreases or the end of the array is reached.
@@
We denote $LCP_{max}$ as the value stored in the root node of the AVL tree on the top of the stack. When a decrease in the $\mathsf{LCP}$ value (or end of array) is encountered, the algorithm pops a collection of AVL trees from the stack, such that all suffixes represented by an AVL tree key have their longest common prefix equal $LCP_{max}$. This collection of AVL trees, denoted as $\mathcal{L}_{suffixSets}$, is then merged into a single AVL tree using the Union operation presented in [12], resulting in $\mathcal{T}_{suffixes}=\bigcup\mathcal{L}_{suffixSets}$. The Union operation also computes a list of ordered pairs representing maximal right-closed substrings via the $ChangeList(\mathcal{L}_{suffixSets},\mathcal{T}_{suffixes})$ operation defined as follows:
$ChangeList(\mathcal{P},\mathcal{P}^{\prime})=\{(x,next_{\mathcal{P}^{\prime}}(x)):next_{\mathcal{P}}(x)\neq next_{\mathcal{P}^{\prime}}(x)\}.$ (5)
@@
where $\mathcal{P}$ is a partition of set $\mathcal{P}^{\prime}$ and for a set $X$ of integers and $x\in X$, $next_{X}(x)=\min\{y\in X\mid y>x\}$. Each pair $(x,y)\in ChangeList(\mathcal{L}_{suffixSets},\mathcal{T}_{suffixes})$ identifies a maximal right-closed substring $r_{j}=w[x..y+LCP_{max}-1]$, where $LCP_{max}$ is the length of its longest border. Hence, the pair $(y+LCP_{max}-x,LCP_{max})$ is added to $\mathcal{MRC}[x]$. All such pairs from the $ChangeList$ are added to the $\mathcal{MRC}$. The algorithm then assigns to the root node of the AVL tree, $\mathcal{T}_{suffixes}$, the $LCP$ value of the AVL tree at the top of the stack and pushes it onto the stack, only if the stack is not empty–thus maintaining the lexicographic ordering of the set of suffixes–and continues scanning the $\mathsf{LCP}$ array.
@@
The $ChangeList(\mathcal{L}_{suffixSets},\mathcal{T}_{suffixes})$ operation then returns every pair of consecutive occurrences of substrings in the $\mathcal{L}_{suffixSets}$. These consecutive occurrences form the border of the maximal right-closed substring and are therefore added to the $\mathcal{MRC}$ array. Finally, the simple linear scan adds maximal right-closed substrings of length one. The algorithm clearly computes the maximal right-closed substrings in strictly decreasing order of length as required in the $\mathcal{MRC}$ array. Thus, we get Theorem 4.1.
We note that the algorithm in the worst case performs $24n-$ 2 stack operations for a string $w=\lambda^{n}$ and in the best case performs $2n$ stack operations for a string with all distinct characters. See Table 1 for an example in the Appendix.
@@
1:$i\leftarrow 0$, $\mathcal{MRC}\leftarrow\{\emptyset\}^{n}$, $stack\leftarrow\emptyset$
@@
13:         $\mathcal{T}_{suffixes}=\bigcup\mathcal{L}_{suffixSets}$triangleright $\bigcup\mathcal{L}_{suffixSets}$ returns an AVL tree
14:         foreach $(x,y)\in ChangeList(\mathcal{L}_{suffixSets},\mathcal{T}_{suffixes})$ do $\triangleright$ Equation 5
@@
Algorithm 1 correctly computes the $\mathcal{MRC}$ array of a string $w[1..n]$ using $\mathsf{SA}$ and $\mathsf{LCP}$.
Theorem 4.2
Algorithm 1 computes the $\mathcal{MRC}$ array of a string $w[1..n]$ using $\mathsf{SA}$ and $\mathsf{LCP}$ in $\mathcal{O}(n\log n)$ time.
@@
We prove the total time complexity of Algorithm 1 in three parts. First, for stack operations, we introduce a potential function $\Phi(S)=|S|\geq 0$, where $|S|$ is the number of elements in the stack. The actual cost of each push and pop operation is 1. Since we push all $n$ single node AVL trees corresponding to each suffix onto the stack, the amortized cost for each push operation is $c^{\prime}_{push}=1+\Delta\Phi(S)=2$, resulting in a total amortized cost of $2n$. Additionally, in the algorithm, we pop at least two AVL trees and push their Union back onto the stack, and because the amortized cost of this sequence of pops and push in the worse case is $c^{\prime}_{pops-push}=3+\Delta\Phi(S)=2$. The maximum number of such operations is $n-1$, leading to a total amortized cost of $2(n-1)$. Therefore, the total cost of all stack operations in the worst case is $24n-$ 2, which is $\mathcal{O}(n)$. Second, by Lemma 4, the time complexity of the sequence of Union operations is $\mathcal{O}(n\log n)$. Finally, the single-characters that are maximal right-closed are computed by a linear scan on $w$, taking $\mathcal{O}(n)$ time. Therefore, the total time complexity is $\mathcal{O}(n\log n)$. ∎
Algorithm to Compute $\mathcal{C}(w)$: Algorithm 2 computes the compact representation $\mathcal{C}(w)$, as in Equation (2), of all closed substrings of a string $w[1..n]$, by simply traversing all the $n$ lists of the $\mathcal{MRC}$ array (see example in Table 2 in the Appendix). Since the $\mathcal{MRC}$ array is of size $\mathcal{O}(n\log n)$ we get Theorem 4.3.
@@
$\mathcal{C}\leftarrow\emptyset$
for $i\leftarrow 1$ to $n$ do
     for $j\leftarrow 1$ to length of $\mathcal{MRC}[i]$ do
         $(\ell r_{j},\ell b_{j})\leftarrow\mathcal{MRC}[i][j]$
         if $j=1$ then
              $\ell p_{j}\leftarrow 1$triangleright Equation (2)
         else
              $(\ell r_{j-1},\ell b_{j-1})\leftarrow\mathcal{MRC}[i][j-1]$
              $\ell p_{j}\leftarrow\ell r_{j}-\ell b_{j}+\ell b_{j-1}+1$triangleright Equation (2)
         insert$(\mathcal{C},(i,\ell p_{j},\ell r_{j}))$
return $\mathcal{C}$
@@
Algorithm 2 correctly computes all $\Theta(n^{2})$ closed substrings of a string $w[1..n]$ in $\mathcal{O}(n\log n)$ time.
Algorithm to Compute All MCSs: Algorithm 3 computes all the MCSs of a string $w[1..n]$ by iterating through the $\mathcal{MRC}$ array (see example in Table 2 in the Appendix). For each index $i$ in $w$ (where $1\leq i\leq n$), the algorithm examines each maximal right-closed substring represented in $\mathcal{MRC}[i]$ and verifies whether the substring is also left-closed using a constant time check $i=1\text{ or }w[i-1]\neq w[i+\ell r-\ell b-1]$. If the condition is satisfied, the substring is added to the MCS list. Since scanning through the $\mathcal{MRC}$ array takes $\mathcal{O}(n\log n)$ time, and each check is performed in constant time, the algorithm computes all MCSs in $\mathcal{O}(n\log n)$ time, establishing Theorem 4.4.
@@
4:         $(\ell r,\ell b)\leftarrow\mathcal{MRC}[i][j]$
5:         if $i=1$ or $w[i-1]\neq w[i+\ell r-\ell b-1]$ then
@@ 5 MCSs in a Fibonacci Word
Let us recall that, a Fibonacci word $f_{n}$ is defined recursively by $f_{0}=0$, $f_{1}=1$, and $f_{n}=f_{n-1}f_{n-2}$ for $n\geq 2$. We denote $F_{n}=|f_{n}|$, where $F_{n}$ satisfies the Fibonacci recurrence relation $F_{n}=F_{n-1}+F_{n-2}$, with $F_{0}=1$ and $F_{1}=1$. We refer to the logical separation between $f_{n-1}$ and $f_{n-2}$ as the boundary.
Bucci et al. [8] study the open and closed prefixes of Fibonacci words and show that the sequence of open and closed prefixes of a Fibonacci word follows the Fibonacci sequence. De Luca et al. [9] explore the sequence of open and closed prefixes of a Sturmian word. In [11], the authors define and find the closed Ziv–Lempel factorization and classify closed prefixes of infinite $m$-bonacci words.
@@
The non-singleton MCSs are either runs or maximal $\alpha$-gapped repeats in which the arm occurs exactly twice [4]. For a Fibonacci word $f_{n}$, let $SM(f_{n})$, $R(f_{n})$, and $GM(f_{n})$ denote the number of singleton MCSs, runs, and gapped-MCSs, respectively. Therefore, the total number of MCSs in $f_{n}$ is given by $M(f_{n})=SM(f_{n})+R(f_{n})+GM(f_{n})$.
@@
Suppose $k+1$ is odd. By inductive hypothesis, $SM(f_{k})=F_{k-2}+F_{k-4}$ and $SM(f_{k-1})=F_{k-3}+F_{k-5}+2$. Since $k$ is even, $f_{k}$ ends with 00 and $f_{k-1}$ begins with $1$, their concatenation results in the string $01$ at the boundary, which does not reduce the count of singleton MCSs in $f_{k+1}$. Thus, we get $SM(f_{k+1})=SM(f_{k})+SM(f_{k-1})=F_{k-1}+F_{k-3}+2$. ∎
@@
To compute $GM(f_{n})$, we count the number of $\alpha$-gapped repeats whose arm occurs exactly twice, as these are, by definition gapped-MCSs. We show that $101$ is the only gapped-MCS in Fibonacci words. Therefore, $GM(f_{n})$ equals the number of occurrences of $101$ in $f_{n}$.
@@
Suppose $uvu$ is a maximal $\alpha$-gapped repeat in $f_{n}$ that is also a suffix of $f_{n}$. Then, the arm $u$ is a Fibonacci word.
@@
For $n\geq 5$, suppose $uvu$ is a maximal $\alpha$-gapped repeat in $f_{n}$ that is also a suffix of $f_{n}$. Then $uvu$ is not a gapped-MCS.
@@
By Lemma 7, the arm of the maximal $\alpha$-gapped repeat is a Fibonacci word $u=f_{k}$. Clearly, $k\neq n$.
If $n$ is even, then the proper suffixes of $f_{n}$ that are Fibonacci words are in the set $\mathcal{S}=\{f_{k}\mid k\mod 2=0\land 0\leq k<n\}$. For any proper suffix $f_{k}\in\mathcal{S}\setminus\{f_{0},f_{2}\}$, $f_{k+2}$ is a suffix of $f_{n}$. By definition, $f_{k+2}=f_{k+1}f_{k}=f_{k}f_{k-1}f_{k}=f_{k}f_{k}f_{k-3}f_{k-2}$, the suffix $f_{k-1}f_{k}=f_{k}f_{k-3}f_{k-2}$ has length $F_{k}+F_{k-1}<2F_{k}$ implying that the last two occurrences of $f_{k}$ overlap. Therefore, any $uvu$ with $f_{k}$ as its arm, will have at least three occurrences of $f_{k}$ making it an open string. For suffix $f_{k}\in\{f_{0},f_{2}\}$ of $f_{n}$, for $n\geq 6$, $f_{6}=1011010110110$ is a suffix of $f_{n}$, and the last two occurrences of $f_{k}$ are left-extendible, and so, it is not maximal. Moreover, any longer $uvu$ with the arm $f_{k}$ will have at least three occurrences of $f_{k}$ making it an open string.
@@
For $n\geq 3$, suppose $uvu$ is a maximal $\alpha$-gapped repeat in $f_{n}$ that is not a suffix of $f_{n}$. Then the arm $u$ belongs to the set $\mathcal{A}_{n}=\{p_{3},p_{4},\ldots,p_{n-2}\}$.
@@
For $n\geq 4$, all the borders of $p_{n}$ are in the set $\mathcal{B}_{n}=\{p_{3},p_{4},\ldots,p_{n-1}\}$.
@@
For $n\geq 5$, all covers of $p_{n}$ belong to the set $\mathcal{C}_{n}=\{p_{4},p_{5},\ldots,p_{n}\}$.
@@
By Lemma 11, for $n\geq 5$, $p_{n-1}$ is a cover of $p_{n}$. By Lemma 10, $p_{n-1}$ is the longest border of $p_{n}$ and therefore the longest proper cover of $p_{n}$. Similarly, $p_{n-2}$ is the longest proper cover of $p_{n-1}$, $p_{n-3}$ is the longest proper cover of $p_{n-2}$, and so on. Therefore, by Lemma 12, it follows that all covers of $p_{n}$ belong to the set $\mathcal{C}_{n}=\{p_{4},p_{5},\ldots,p_{n}\}$ for all $n\geq 5$. ∎
@@
For $n\geq 5$, the substring $101$ is the only gapped-MCS in $f_{n}$.
@@
All gapped-MCSs are maximal $\alpha$-gapped repeats. By Lemma 8, for $n\geq 5$, all maximal $\alpha$-gapped repeats that are also suffixes of $f_{n}$ are not gapped-MCSs.
By Lemma 9, for $n\geq 3$, if $uvu$ is a maximal $\alpha$-gapped repeat in $f_{n}$ that is not a suffix of $f_{n}$, then the arm $u$ belongs to the set $\mathcal{A}_{n}=\{p_{3},p_{4},\ldots,p_{n-2}\}$. By Lemma 13, all covers of $p_{n}$ belong to the set $\mathcal{C}_{n}=\{p_{4},p_{5},\ldots,p_{n}\}$.
Suppose there exists a maximal $\alpha$-gapped $w=uvu$, where $u\in\mathcal{A}_{n}\setminus{p_{3}}$, that is a gapped-MCS. Assume $n$ is odd. Then, $f_{n}=p_{n}\delta_{n}$, where $\delta_{n}=01$, and $f_{n}[F_{n}-1]=0$. Since $u\in\mathcal{A}_{n}\setminus{p_{3}}$, it ends with $1$. Clearly, $u$ does not occur ending at position $F_{n}-1$. Now suppose $n$ is even. By a similar reasoning, $u$ does not occur ending at $F_{n}-1$ since it ends with $01$ and $f_{n}[F_{n}-2]=f_{n}[F_{n}-1]=1$ (as $f_{4}=10110$ is a suffix of $f_{n}$). Hence, $u$ ends within $p_{n}$. Since $u$ is a cover of $p_{n}$, $w=uvu$ will contain at least three occurrences of $u$ making it an open string. Therefore, there is no maximal $\alpha$-gapped repeat of the form $w=uvu$, where $u\in\mathcal{A}_{n}\setminus{p_{3}}$, that is a gapped MCS.
Now we consider the final case, where $u=p_{3}=1$. In this case, we get $101$ as the only gapped-MCS. Any longer maximal $\alpha$-gapped repeat will have at least three occurrences of $u=p_{3}=1$, making it an open string—a contradiction. ∎

### v2

@@ 4 Algorithms for Closed Substrings and MCSs
In this section, we introduce the primary data structure –the $\mathcal{MRC}$ array– which is used to compute all closed substrings in a compact representation and MCSs in $\mathcal{O}(n\log n)$ time. See Table 1 for an example of the $\mathcal{MRC}$ array for the string $w[1..11]=mississippi$ in the Appendix.
Definition 1($\mathcal{MRC}$ Array [^2] This definition uses the notation introduced in Section 3.)
@@
Given a string $w[1..n]$, the maximal right-closed array ($\mathcal{MRC}$) is an array of size $n$, where $\mathcal{MRC}[i]$, $1\leq i\leq n$, stores the list of ordered pairs $(|r_{j}|,|b_{j}|)$, where $j=K_{i}$ to $1$.
@@
Algorithm 1 computes the $\mathcal{MRC}$ array by first identifying all the maximal right-closed substrings of length greater than one. It uses a stack to store AVL trees created during its execution. The algorithm scans the $\mathsf{LCP}$ array from left to right. Beginning with $\mathsf{LCP}[1]=0$, and for increasing $\mathsf{LCP}[i]$ values, the algorithm creates a single node (root) AVL tree with $key=\mathsf{SA}[i]$ and $value=\mathsf{LCP}[i]$ and pushes it onto the stack, until either the $LCP$ value decreases or the end of the array is reached.
@@
We denote by $LCP_{max}$ the value stored in the root node of the AVL tree on the top of the stack. When a decrease in the $LCP$ value (or the end of the array) is encountered, the algorithm pops a collection of AVL trees from the stack, denoted as $\mathcal{L}_{suffixSets}$, such that all suffixes represented by an AVL tree key share the same longest common prefix equal to $LCP_{max}$. Since $LCP_{max}$ is the maximum $LCP$ value of all the AVL trees on the stack, popping this collection, $\mathcal{L}_{suffixSets}$, preserves the strictly decreasing order of ordered pairs in the $\mathcal{MRC}$ array. This collection is then merged into a single AVL tree using the Union operation presented in [16] [^3] The Union operation, defined in [16], builds upon foundational methods from [8, 9]., resulting in $\mathcal{T}_{suffixes}=\bigcup\mathcal{L}_{suffixSets}$. The Union operation also computes a list of ordered pairs representing maximal right-closed substrings via the $ChangeList(\mathcal{L}_{suffixSets},\mathcal{T}_{suffixes})$ operation defined as follows:
$ChangeList(\mathcal{P},\mathcal{P}^{\prime})=\{(x,next_{\mathcal{P}^{\prime}}(x)):next_{\mathcal{P}}(x)\neq next_{\mathcal{P}^{\prime}}(x)\},$ (5)
@@
where $\mathcal{P}$ is a partition of the set $\mathcal{P}^{\prime}$, and $next_{X}(x)=\min\{y\in X\mid y>x\}$, $x\in X$ and $X\subseteq\mathbb{Z}^{+}$. In particular, if $x\in P_{i}$ for some $P_{i}\in\mathcal{P}$, then $next_{\mathcal{P}}(x)=next_{P_{i}}(x)$. Each pair $(x,y)\in ChangeList(\mathcal{L}_{suffixSets},\mathcal{T}_{suffixes})$ identifies a maximal right-closed substring $r_{j}=w[x..y+LCP_{max}-1]$, where $LCP_{max}$ is the length of its longest border. Hence, the pair $(y+LCP_{max}-x,LCP_{max})$ is added to $\mathcal{MRC}[x]$. All such pairs from the $ChangeList$ are added to the $\mathcal{MRC}$ array. The algorithm then assigns to the root node of the AVL tree, $\mathcal{T}_{suffixes}$, the $LCP$ value of the AVL tree at the top of the stack and pushes it onto the stack, only if the stack is not empty–thus maintaining the lexicographic ordering of the set of suffixes–and continues scanning the $\mathsf{LCP}$ array.
@@
The $ChangeList(\mathcal{L}_{suffixSets},\mathcal{T}_{suffixes})$ operation then returns every pair of consecutive occurrences of substrings in the $\mathcal{L}_{suffixSets}$. These consecutive occurrences form the border of the maximal right-closed substring and are therefore added to the $\mathcal{MRC}$ array. Finally, the simple linear scan adds maximal right-closed substrings of length one. The algorithm clearly computes the maximal right-closed substrings in strictly decreasing order of length as required in the $\mathcal{MRC}$ array. Thus, the correctness of Algorithm 1 follows.
We note that the algorithm in the worst case performs 4n24n-2 stack operations for a string $w=\lambda^{n}$ and in the best case performs $2n$ stack operations for a string with all distinct characters.
@@
[…]
@@
13:   $\mathcal{T}_{suffixes}=\bigcup\mathcal{L}_{suffixSets}$ $\triangleright$ $\bigcup\mathcal{L}_{suffixSets}$ returns an AVL tree
14:   foreach $(x,y)\in ChangeList(\mathcal{L}_{suffixSets},\mathcal{T}_{suffixes})$ do $\triangleright$ Equation (5)
@@
Algorithm 1 correctly computes the $\mathcal{MRC}$ array of a string $w[1..n]$ using $\mathsf{SA}$ and $\mathsf{LCP}$ in $\mathcal{O}(n\log n)$ time.
@@
We prove the total time complexity of Algorithm 1 in three parts. First, for stack operations, we introduce a potential function $\Phi(S)=|S|\geq 0$, where $|S|$ is the number of elements in the stack. The actual cost of each push and pop operation is 1. Since we push all $n$ single node AVL trees corresponding to each suffix onto the stack, the amortized cost for each push operation is $c^{\prime}_{push}=1+\Delta\Phi(S)=2$, resulting in a total amortized cost of $2n$. Additionally, in the algorithm, we pop at least two AVL trees and push their Union back onto the stack, and because the amortized cost of this sequence of pops and push in the worse case is $c^{\prime}_{pops-push}=3+\Delta\Phi(S)=2$. The maximum number of such operations is n1n-1, leading to a total amortized cost of 2(n1)2(n-1). Therefore, the total cost of all stack operations in the worst case is 4n24n-2, which is $\mathcal{O}(n)$. Second, by Lemma 4, the time complexity of the sequence of Union operations is $\mathcal{O}(n\log n)$. Finally, the single-characters that are maximal right-closed are computed by a linear scan on $w$, taking $\mathcal{O}(n)$ time. Therefore, the total time complexity is $\mathcal{O}(n\log n)$. ∎
Algorithm to Compute $\mathcal{C}(w)$: Algorithm 2 computes the compact representation $\mathcal{C}(w)$, as in Equation (2), of all closed substrings of a string $w[1..n]$, by simply traversing all $n$ lists in the $\mathcal{MRC}$ array (see example in Table 2 in the Appendix). Since the $\mathcal{MRC}$ array is of size $\mathcal{O}(n\log n)$ we get Theorem 4.2.
@@
[…]
2:for $i\leftarrow 1$ to $n$ do
3:  for $j\leftarrow 1$ to length of $\mathcal{MRC}[i]$ do
4:   $(\ell r_{j},\ell b_{j})\leftarrow\mathcal{MRC}[i][j]$ $\triangleright$ $\ell r_{j}=|r_{j}|,\ell b_{j}=|b_{j}|,\ell p_{j}=|p_{j}|$
5:   if $j=1$ then
6:     $\ell p_{j}\leftarrow 1$ $\triangleright$ Equation (2)
7:   else
8:     $(\ell r_{j-1},\ell b_{j-1})\leftarrow\mathcal{MRC}[i][j-1]$
9:     $\ell p_{j}\leftarrow\ell r_{j}-\ell b_{j}+\ell b_{j-1}+1$ $\triangleright$ Equation (2)
[…]
11:return $\mathcal{C}$
@@
Algorithm 2 correctly computes all $\mathcal{O}(n^{2})$ closed substrings of a string $w[1..n]$ in a compact representation in $\mathcal{O}(n\log n)$ time.
Algorithm to Compute All MCSs: Algorithm 3 computes all MCSs of a string $w[1..n]$ by iterating through the $\mathcal{MRC}$ array (see example in Table 2 in the Appendix). For each index $i$ in $w$ (where $1\leq i\leq n$), the algorithm examines each maximal right-closed substring represented in $\mathcal{MRC}[i]$ and verifies whether the substring is also maximal left-closed using a constant time check $i=1$ or $(i>1\text{ and }w[i-1]\neq w[i+|r|-|b|-1])$. If the condition is satisfied, the substring is added to the MCS list. Since scanning through the $\mathcal{MRC}$ array takes $\mathcal{O}(n\log n)$ time, and each check is performed in constant time, the algorithm computes all MCSs in $\mathcal{O}(n\log n)$ time, establishing Theorem 4.3.
@@
4:   $(\ell r,\ell b)\leftarrow\mathcal{MRC}[i][j]$ $\triangleright$ $\ell r=|r|,\ell b=|b|$
5:   if $i=1$ or $(i>1\textbf{ and }w[i-1]\neq w[i+\ell r-\ell b-1])$ then
@@ 5 MCSs in a Fibonacci Word
Let us recall that a Fibonacci word $f_{n}$ is defined recursively by $f_{0}=0$, $f_{1}=1$, and $f_{n}=f_{n-1}f_{n-2}$ for $n\geq 2$. We denote $F_{n}=|f_{n}|$, where $F_{n}$ satisfies the Fibonacci recurrence relation $F_{n}=F_{n-1}+F_{n-2}$, with $F_{0}=1$ and $F_{1}=1$. We refer to the logical separation between $f_{n-1}$ and $f_{n-2}$, in the string $f_{n}$, as the boundary.
Bucci et al. [10] study the open and closed prefixes of Fibonacci words and show that the sequence of open and closed prefixes of a Fibonacci word follows the Fibonacci sequence. De Luca et al. [11] explore the sequence of open and closed prefixes of a Sturmian word. In [14], the authors define and find the closed Ziv–Lempel factorization and classify closed prefixes of infinite $m$-bonacci words. In this section, we extend previous work on closed substrings in Fibonacci words by deriving a formula to compute the number of MCSs they contain.
@@
The non-singleton MCSs are either runs or maximal gapped repeats in which the arm occurs exactly twice [4]. For a Fibonacci word $f_{n}$, let $SM(f_{n})$, $R(f_{n})$, and $GM(f_{n})$ denote the number of singleton MCSs, runs, and gapped-MCSs, respectively. Therefore, the total number of MCSs in $f_{n}$ is given by $M(f_{n})=SM(f_{n})+R(f_{n})+GM(f_{n})$.
@@
Suppose $k+1$ is odd. By inductive hypothesis, $SM(f_{k})=F_{k-2}+F_{k-4}$ and $SM(f_{k-1})=F_{k-3}+F_{k-5}+2$. Since $k$ is even, $f_{k}$ ends with 0 and $f_{k-1}$ begins with $1$, their concatenation results in the string $01$ at the boundary, which does not reduce the count of singleton MCSs in $f_{k+1}$. Thus, we get $SM(f_{k+1})=SM(f_{k})+SM(f_{k-1})=F_{k-1}+F_{k-3}+2$. ∎
@@
To compute $GM(f_{n})$, we count the number of maximal gapped repeats whose arms occur exactly twice, as these are, by definition gapped-MCSs. In Theorem 5.1 we first show that the only gapped-MCSs in a Fibonacci word are occurrences of the substring $101$ that are also maximal gapped repeats. Then, we count all such occurrences in $f_{n}$.
@@
Suppose uvuuvu is a maximal gapped repeat in $f_{n}$ that is also a suffix of $f_{n}$. Then, the arm $u$ is a Fibonacci word.
@@
For $n\geq 5$, suppose uvuuvu is a maximal gapped repeat in $f_{n}$ that is also a suffix of $f_{n}$. Then uvuuvu is not a gapped-MCS.
@@
By Lemma 7, the arm of the maximal gapped repeat is a Fibonacci word $u=f_{k}$. Clearly, $k\neq n$.
If $n$ is even, then the proper suffixes of $f_{n}$ that are Fibonacci words are in the set $\mathcal{S}=\{f_{k}\mid k\mod 2=0\land 0\leq k<n\}$. For any proper suffix $f_{k}\in\mathcal{S}\setminus\{f_{0},f_{2}\}$, $f_{k+2}$ is a suffix of $f_{n}$. By definition, $f_{k+2}=f_{k+1}f_{k}=f_{k}f_{k-1}f_{k}=f_{k}f_{k}f_{k-3}f_{k-2}$, the suffix $f_{k-1}f_{k}=f_{k}f_{k-3}f_{k-2}$ has length $F_{k}+F_{k-1}<2F_{k}$ implying that the last two occurrences of $f_{k}$ overlap. Therefore, any uvuuvu with $f_{k}$ as its arm, will have at least three occurrences of $f_{k}$ making it an open string. For suffix $f_{k}\in\{f_{0},f_{2}\}$ of $f_{n}$, for $n\geq 6$, $f_{6}=1011010110110$ is a suffix of $f_{n}$, and the last two occurrences of $f_{k}$ are left-extendible, and so, it is not maximal. Moreover, any longer uvuuvu with the arm $f_{k}$ will have at least three occurrences of $f_{k}$ making it an open string.
@@
For $n\geq 3$, suppose uvuuvu is a maximal gapped repeat in $f_{n}$ that is not a suffix of $f_{n}$. Then the arm $u$ belongs to the set $\mathcal{A}_{n}=\{p_{3},p_{4},\ldots,p_{n-2}\}$.
@@
For $n\geq 4$, all the borders of $p_{n}$ form the set $\mathcal{B}_{n}=\{p_{3},p_{4},\ldots,p_{n-1}\}$.
@@
For $n\geq 5$, all covers of $p_{n}$ form the set $\mathcal{C}_{n}=\{p_{4},p_{5},\ldots,p_{n}\}$.
@@
By Lemma 11, for $n\geq 5$, $p_{n-1}$ is the longest proper cover of $p_{n}$. Similarly, $p_{n-2}$ is the longest proper cover of $p_{n-1}$, $p_{n-3}$ is the longest proper cover of $p_{n-2}$, and so on. Therefore, by Lemma 12, it follows that all covers of $p_{n}$ belong to the set $\mathcal{C}_{n}=\{p_{4},p_{5},\ldots,p_{n}\}$ for all $n\geq 5$. ∎
@@
For $n\geq 5$, every gapped-MCS in a Fibonacci word $f_{n}$ corresponds to an occurrence of the substring $101$ that is also a maximal gapped repeat.
@@
All gapped-MCSs are maximal gapped repeats. By Lemma 8, for $n\geq 5$, all maximal gapped repeats that are also suffixes of $f_{n}$ are not gapped-MCSs.
By Lemma 9, for $n\geq 3$, if uvuuvu is a maximal gapped repeat in $f_{n}$ that is not a suffix of $f_{n}$, then the arm $u$ belongs to the set $\mathcal{A}_{n}=\{p_{3},p_{4},\ldots,p_{n-2}\}$. By Lemma 13, all covers of $p_{n}$ belong to the set $\mathcal{C}_{n}=\{p_{4},p_{5},\ldots,p_{n}\}$.
Suppose there exists a maximal gapped $w=uvu$, where $u\in\mathcal{A}_{n}\setminus{p_{3}}$, that is a gapped-MCS. Assume $n$ is odd. Then, $f_{n}=p_{n}\delta_{n}$, where $\delta_{n}=01$, and $f_{n}[F_{n}-1]=0$. Since $u\in\mathcal{A}_{n}\setminus{p_{3}}$, it ends with $1$. Clearly, $u$ does not occur ending at position $F_{n}-1$. Now suppose $n$ is even. By a similar reasoning, $u$ does not occur ending at $F_{n}-1$ since it ends with $01$ and $f_{n}[F_{n}-2]=f_{n}[F_{n}-1]=1$ (as $f_{4}=10110$ is a suffix of $f_{n}$). Hence, $u$ ends within $p_{n}$. Since $u$ is a cover of $p_{n}$, $w=uvu$ will contain at least three occurrences of $u$ making it an open string. Therefore, there is no maximal gapped repeat of the form $w=uvu$, where $u\in\mathcal{A}_{n}\setminus{p_{3}}$, that is a gapped MCS.
Now we consider the final case, where $u=p_{3}=1$. In this case, we get $101$ as the only gapped-MCS. Any longer maximal gapped repeat will have at least three occurrences of $u=p_{3}=1$, making it an open string — a contradiction. ∎